# On cases where Litt’s game is fair

### v1

@@ Abstract
A fair coin is flipped $n$ times, and two finite sequences of heads and tails (words) $A$ and $B$ of the same length are given. Each time the word $A$ appears in the sequence of coin flips, Alice gets a point, and each time the word $B$ appears, Bob gets a point. Who is more likely to win? This puzzle is a slight extension of Litt’s game [4] that recently set Twitter abuzz. We show that Litt’s game is fair for any value of $n$ and any two words that have the same auto-correlation structure by building up a bijection that exchanges Bob and Alice scores; the fact that the inter-correlation does not come into play in this case may come up as a surprise.
@@ 1 Introduction and main results
    for large $n$, it holds:
    $2\cdot\big{(}\mathbb{P}_{n}(\mbox{Bob wins})-\mathbb{P}_{n}(\mbox{Alice wins})\big{)}\sim\mathbb{P}_{n}(\mbox{Tie})\sim\frac{1}{\sqrt{\pi n}}.$
Several elements of proofs for the first fact quickly emerged on Twitter [8, 9, 11, 12], then more formal proofs of both facts appeared on the arXiv [1, 10]. Nica [5] recorded a YouTube video to introduce the problem to a wider audience. Zeilberger online journal [13] maintains a list of contributions and questions on the problem; among the proofs given, we would like to advertise an early probabilistic proof: in [8], Ramesh builds a ”fair upper-bound” for the score difference of Alice and Bob that consists in a delayed simple random walk; the score difference of Alice and Bob is a deterministic function of this walk, at distance at most 1 below it, see right after the bibliography for more details. A rigorous proof of both facts 1 and 2 is then at an easy reach. Another by-product of his approach is that giving an initial advantage of only one point to Alice reverses the statement of the first fact in a strong sense: for any size $n\geqslant 0$, Alice then has probability ¿1/2 of winning; in particular, Alice is more likely to win than Bob.
Considering a contest between HT and TH is also possible, but arguably less interesting: flipping the sequence of tosses, or reading the sequence in the reverse order exchanges Alice and Bob points, which results in a fair game. If you found this second contest boring, beware: the purpose of this short note is a generalization of it - the symmetry being slightly more hidden though…
For $\ell\geqslant 0$, a word of length $\ell$ is a sequence $A=(a_{1},\ldots,a_{\ell})$ in $\{0,1\}^{\ell}$. For $X_{n}:=(\varepsilon_{k})_{1\leqslant k\leqslant n}$ a finite sequence in $\{0,1\}^{n}$, we denote by
@@
the number of occurrences of the word $A$ in the sequence $X_{n}$. Assume now that $X_{n}$ is an i.i.d. sequence of Bernoulli random variables with parameter $1/2$ and let $A$ and $B$ be two words of size $\ell$. In the generalized Litt’s game, Alice wins if $N_{A}(X_{n})>N_{B}(X_{n})$, Bob wins if $N_{B}(X_{n})>N_{A}(X_{n})$, and this is a tie otherwise. A key quantity encoding the intersections of $A$ and $B$ is the correlation of $A$ and $B$, which may be represented as a subset of integers or simply as a number (the base 2 expansion of the subset):
@@
To be more specific, we shall call inter-correlation the correlation of two distinct words (beware the order matters), and auto-correlation the correlation of a word with itself. Our main result in this note is that Litt’s game is fair for words $A$ and $B$ with the same auto-correlation, whatever their inter-correlations may be.
@@
Let $A$, $B$ two words of length $\ell$ such that $[A|A]=[B|B]$. Let $X_{n}=(\varepsilon_{i})_{1\leqslant i\leqslant n}$ be an i.i.d. sequence with law $\mathcal{B}(1/2)$. Then for each $n\geqslant 1$, $(N_{A}(X_{n}),N_{B}(X_{n}))$ and $(N_{B}(X_{n}),N_{A}(X_{n}))$ have the same distribution. In particular, for any $n\geqslant 1$
@@
To illustrate this result, let us pick for instance $A=HHTHTH$ and $B=HTTTHH$. These words have the same auto-correlation so Litt’s game is perfectly fair yet there does not seem to be an apparent symmetry readily implying this fact.
It is possible to estimate this quantity, or equivalently the probability of ties $\mathbb{P}_{n}(\text{Ties})$, using a Local Central Limit theorem for sums of weakly dependent random variables. We do not venture into this for the following reason: in a recent breakthrough announced online [6], Nica (together with Janson) pointed to a generic method based on Edgeworth expansions to tackle the case of words $A$ and $B$ with possibly distinct auto-correlations; the sketch of proof, together with some basic moment computations, hints to the following (yet still conjectural at the moment of writing) asymptotic estimates: for $A\neq B$, as $n$ gets large, the following asymptotics:
@@
$\mathbb{P}_{n}(\mbox{Ties})$ $=\frac{2^{\ell}}{\sqrt{2^{\ell}+[A|A]+[B|B]-[A|B]-[B|A]}}\cdot\frac{1}{2\sqrt{\pi n}}+o(\frac{1}{\sqrt{n}}\Big{)}$
hold as soon as the denominators on both RHS are non null (the four111$(TH^{\ell-1},H^{\ell-1}T)$ and its siblings, for which $N_{A}(X_{n})-N_{B}(X_{n})$ belongs to $\{-1,0,1\}$. pair of words giving a null denominator being associated with degenerated cases). In case $[A|A]=[B|B]$, Theorem 1 refines on the first formula by stating that the RHS is, in fact, identically null for all $n$.
A key feature of the first formula is that for $\mathbb{P}_{n}(\mbox{Bob wins})-\mathbb{P}_{n}(\mbox{Alice wins})$ has the sign of $[A|A]-[B|B]$. In light of this observation together with some extensive numerical computations, we are lead to state the following conjecture for a fixed length $n$ of the underlying word. We use $\Delta$ to denote the symmetric difference of two sets.
@@
For each222for smaller values of $n$, the game is fair. $n\geqslant 2\ell-\max\{\hbox{{Cor}}(A)\Delta\hbox{{Cor}}(B)\}$,
$\mathbb{P}_{n}(\emph{Bob wins})-\mathbb{P}_{n}(\emph{Alice wins})\text{ has the sign of }[A|A]-[B|B],$
meaning these quantities are either both positive, null, or negative.
Our main result, Theorem 1, answers the conjecture in case the two words have the same auto-correlation, $[A|A]=[B|B]$. This case would also be, according to our conjecture, the only case where Litt’s game is fair. Last, the inequality $|[A|A]-[B|B]|\leqslant 2^{\ell}-2$ holds for every pair of words $A$,$B$, hence the probability of ties is asymptotically always larger than the absolute value of the difference of the win probabilities for Alice and Bob.
We should also point that the topic of pattern matching and overlaps has been the subject of many investigations, starting with the celebrated non-transitive Penney Ante game named after Penney [7], who asked about the first word to appear (in between two words) in a sequence of fair coins, see [2, 3] and the references therein.
@@ 2 Proof of Theorem 1
Fix any two words $A$ and $B$ with length $\ell$ and the same auto-correlation. To prove that $(N_{A}(X_{n}),N_{B}(X_{n}))$ has the same law as $(N_{B}(X_{n}),N_{A}(X_{n}))$, we prove the existence of a bijection $\phi$ from $\{0,1\}^{n}$ to $\{0,1\}^{n}$ such that, for any sequence $X_{n}\in\{0,1\}^{n}$, $(N_{A}(X_{n}),N_{B}(X_{n}))=(N_{B}(\phi(X_{n})),N_{A}(\phi(X_{n})))$.
We introduce some additional notation. If $C,D$ are two words, $CD$ will be the concatenation of $C$ and $D$. Besides, for two words $C,D$ of length $\ell$ and $m\in\hbox{{Cor}}(C,D)$, we denote $C^{m}D$ the word of length $2\ell-m$ beginning by $C$ and ending by $D$. For example, if $C=\underline{100101}$, $D=\overline{010111}$, we have $C^{2}D=\underline{1001}\underline{\overline{\mathbf{01}}}\overline{0111}$. We extend this notation to $k$ words $C_{1},\ldots,C_{k}$ of length $\ell$ and $m_{1},\ldots,m_{k}\geqslant 1$ in the obvious way: if $m_{i}\in\hbox{{Cor}}(C_{i},C_{i+1})$ for each $i$, we define the word of length $k\ell-\sum m_{i}$
@@
Let $A,B$ be two words with the same auto-correlation. Let $\phi:\mathcal{E}(A,B)\to\mathcal{E}(A,B)$ define in the following way. If $Y:=C_{1}\mbox{}^{m_{1}}C_{2}\mbox{}^{m_{2}}\ldots C_{k-1}\mbox{}^{m_{k-1}}C_{k}$ with $C_{i}\in\{A,B\}$, we set

### v2

@@ Abstract
A fair coin is flipped $n$ times, and two finite sequences of heads and tails with the same length are given, say $A$ and $B$. Each time the word $A$ appears in the sequence of fair coin flips, Alice gets a point, and each time the word $B$ appears, Bob gets a point. Who is more likely to win? This puzzle is a slight extension of Litt’s game [8] that recently set Twitter abuzz. We show that Litt’s game is fair for any value of $n$ and any two words that have the same auto-correlation structure by building up a bijection that exchanges Bob and Alice scores; the fact that the inter-correlations do not come into play in this case may come up as a surprise.
@@ 1 Introduction and main results
    for $n$ large, it holds:
    $2\times\big{(}\mathbb{P}_{n}(\mbox{Bob wins})-\mathbb{P}_{n}(\mbox{Alice wins})\big{)}\sim\mathbb{P}_{n}(\mbox{Tie})\sim\frac{1}{\sqrt{\pi n}}.$
Several elements of proofs for the first fact quickly emerged on Twitter [13, 14, 17, 18], then more formal proofs of both facts appeared on the arXiv [3, 16]. Nica [10] recorded a YouTube video to introduce the problem to a wider audience. Zeilberger online journal [19] maintains a list of contributions and questions on the problem; among the proofs given, we would like to advertise an early probabilistic proof : in [13], Ramesh builds a “fair majorant” for the score difference of Alice and Bob that consists in a delayed simple random walk; the score difference of Alice and Bob is a deterministic function of this walk, at distance at most 1 below it, see right after the bibliography for more details. A rigorous proof of both facts 1 and 2 is then at an easy reach. Another by-product of his approach is that giving Alice an initial advantage of only one point reverses the statement of fact 1 in a strong sense: for any size $n\geqslant 0$, Alice then has probability strictly greater than $1/2$ of winning. In particular, Alice is more likely to win than Bob. This approach has been detailed in Grimmett [5], who recently rediscovered the arguments of Ramesh.
Considering a contest between HT and TH is also possible, but arguably less interesting: flipping the sequence of tosses, or reading the sequence in the reverse order exchanges Alice and Bob points, which results in a fair game. Our aim in this note is to uncover more hidden symmetries within the words ensuring the game is again tied up.
For $\ell\geqslant 0$, a word of length $\ell$ is a sequence $A=(a_{1},\ldots,a_{\ell})$ in $\{\text{H},\text{T}\}^{\ell}$. For $X_{n}:=(\varepsilon_{k})_{1\leqslant k\leqslant n}$ a finite sequence in $\{\text{H},\text{T}\}^{n}$, we denote by
@@
the number of occurrences of the word $A$ in the sequence $X_{n}$. Let $A$ and $B$ be two words of length $\ell$. In the generalized Litt’s game, Alice wins if $N_{A}(X_{n})>N_{B}(X_{n})$, Bob wins if $N_{B}(X_{n})>N_{A}(X_{n})$, and this is a Tie otherwise. A key quantity encoding the intersections of $A$ and $B$ is the correlation of $A$ and $B$, which may be represented as a subset of integers or simply as a number (the base 2 expansion of the subset):
@@
To be more specific, we shall call inter-correlation the correlation of two distinct words (beware the order matters), and auto-correlation the correlation of a word with itself. Our main result in this note is that Litt’s game, which is played with a uniform $X_{n}$, is fair for words $A$ and $B$ with the same auto-correlation, regardless of their inter-correlation.
@@
[…]
For instance, the words $A=$ HHTHTH and $B=$ HTTTHH have the same auto-correlation 2 hence Litt’s game played with $A$ and $B$ is fair despite the apparent lack of symmetry between these words.
It is then possible to estimate the quantity in the last equality, or equivalently the probability of a tie $\mathbb{P}_{n}(\text{Tie})$, using a Local Central Limit theorem for sums of weakly dependent random variables. We do not venture into this for the following reason : In a recent breakthrough announced online [11], Nica (together with Janson) pointed to a generic method based on Edgeworth expansions to tackle the case of words $A$ and $B$ with possibly distinct auto-correlations; the sketch of proof, together with some basic moment computations, hints to the following asymptotic estimates: for $A\neq B$, as $n$ gets large, the following asymptotics:
@@
$\mathbb{P}_{n}(\mbox{Tie})$ $=\frac{2^{\ell}}{\sqrt{2^{\ell}+[A|A]+[B|B]-[A|B]-[B|A]}}\cdot\frac{1}{2\sqrt{\pi n}}+o(\frac{1}{\sqrt{n}}\Big{)}$
hold as soon as the denominators on both RHS are non null (the four111(THℓ-1,Hℓ-1T) and its siblings, for which $N_{A}(X_{n})-N_{B}(X_{n})$ belongs to $\{-1,0,1\}$. pair of words giving a null denominator are associated with degenerated cases). Shortly after posting a first version of this work, Segert communicated to us a proof, available online [15] as an addendum to his arXiv paper, of a general statement that entails both estimates. In case $[A|A]=[B|B]$, Theorem 1 refines on the first formula by stating that the quantity is, in fact, identically null for all $n$.
A key feature of the first formula is that $\mathbb{P}_{n}(\mbox{Bob wins})-\mathbb{P}_{n}(\mbox{Alice wins})$ has the sign of $[A|A]-[B|B]$. In light of this observation together with some extensive numerical computations222the conjecture has been checked for all words $A$ and $B$ of size $\leqslant 10$ and for $n\leqslant 35$., we are lead to state the following remarkable conjecture for a fixed length $n$ of the underlying word $X_{n}$. We use $\Delta$ to denote the symmetric difference of two sets.
@@
Assume $[A|A]>[B|B]$. Then, for each $n\geqslant n_{0}:=2\ell-\max\{\hbox{{Cor}}(A)\Delta\hbox{{Cor}}(B)\}$,
[…]
Some comments are in order. First, the restriction to $n\geqslant n_{0}$ is necessary for the first inequality only (the most important one indeed), and we have equality for $n<n_{0}$. Also, the last two inequalities are equivalent (and valid for each $n\geqslant 0$), but we like to phrase the conjecture this way, because of the following vivid interpretation of this set of inequalities: even if Alice looses in the standard version of the game (first inequality), if we were to give Alice one additional point 333beware this is distinct from starting the random word $X_{n}$ with an $A$, then her new winning probability would exceed $1/2$ for each $n$ (last inequality); in particular, Alice would be advantaged over Bob at each subsequent time. Conjecture 1 also suggests that the advantage of Bob versus Alice may be put in bijection with a subset of the Tie event, a key feature of the proof by Ramesh indeed. Conjecture 1 is valid for the original Litt’s game (HH versus HT) by Ramesh’s proof. Also, it is in line with the asymptotics above: to see this, notice the last two inequalities in the conjecture are equivalent to
$|\mathbb{P}_{n}(\text{Alice wins})-\mathbb{P}_{n}(\text{Bob wins})|<\mathbb{P}_{n}(\text{Tie}),$
which may be checked from the conjectured asymptotics because of the inequality $|[A|A]-[B|B]|\leqslant 2^{\ell}-2$ that is true for every pair of words $A$,$B$.
We should also point that the topic of pattern matching and overlaps has been the subject of many investigations, starting with the non-transitive Penney Ante game named after Penney [12] (and famously solved by Conway) in which one is asked to compute the probability that $A$ appears before $B$ in a sequence of fair coin flips, see [6, 7] and the references therein.
Our fixed length conjecture is reminiscent of an important fixed length result in the area of pattern matching : the fact that the number of words $X$ of length $n$ avoiding $A$ is larger than the number of words avoiding $B$ if $[A,A]>[B,B]$ for each $n\geqslant n_{0}$ defined in term of $A$ and $B$ as in Conjecture 1, and with equality before. For $n$ large enough (larger than an indefinite constant), the statement has been stated and proved by Guibas and Odlyzko [6] and independently by Blom [1], while the result for $n\geqslant n_{0}$ has been established by Månnson [9] after a series of works on the topic [4, 2].
@@ 2 Proof of Theorem 1
[…]
We introduce some additional notation. If $C,D$ are two words, $CD$ will be the concatenation of $C$ and $D$. Besides, for two words $C,D$ of length $\ell$ and $m\in\hbox{{Cor}}(C,D)$, we denote $C^{m}D$ the word of length $2\ell-m$ beginning by $C$ and ending by $D$. For example, if $C=\underline{\text{HTTHTH}}$, $D=\overline{\text{THTHHH}}$, we have $C^{2}D=\underline{\text{HTTH}}\underline{\overline{\mathbf{\text{TH}}}}\overline{\text{THHH}}$. We extend this notation to $k$ words $C_{1},\ldots,C_{k}$ of length $\ell$ and $m_{1},\ldots,m_{k}\geqslant 1$ in the obvious way: if $m_{i}\in\hbox{{Cor}}(C_{i},C_{i+1})$ for each $i$, we define the word of length $k\ell-\sum m_{i}$
@@
Let $A,B$ be two words with the same auto-correlation, and let $\phi:\mathcal{E}(A,B)\to\mathcal{E}(A,B)$ be defined in the following way. If $Y:=C_{1}\mbox{}^{m_{1}}C_{2}\mbox{}^{m_{2}}\ldots C_{k-1}\mbox{}^{m_{k-1}}C_{k}$ with $C_{i}\in\{A,B\}$, we set